OpenStudy (anonymous):

can someone show me an explained example of probability distribution please?

OpenStudy (zarkon):

\[P(X=x)=p^{x}(1-p)^{1-x}\] \[x\in\{0,1\}\] \[0\le p\le1\]

OpenStudy (anonymous):

what does that all mean?

OpenStudy (zarkon):

it is a bernoulli random variable

OpenStudy (zarkon):

\[P(X=0)=(1-p)\] \[P(X=1)=p\]

OpenStudy (anonymous):

what does everything resemble?

OpenStudy (zarkon):

{0,1} is the sample space p is the probability of a 'success'

OpenStudy (zarkon):

you ever study the binomial distribution?

OpenStudy (anonymous):

honestly.. i can't answer that question.. but if i have then i don't remember it. Its been a few good years since a math course (actually so long i dont remember how to use the graphing calc anymore) but I have to have finite math in order to take my last class to finish my degree

OpenStudy (zarkon):

ic... the Bernoulli distribution is probably the simplest discrete distribution.

OpenStudy (anonymous):

whats the binomial distribution you were talking about?

OpenStudy (zarkon):

a bernoulli dist is a binomial distribution with n=1 the dist is \[P(X=x)={n\choose x}p^x(1-p)^{n-x}\]

OpenStudy (zarkon):

if n=1 you get bernoulli

OpenStudy (anonymous):

lol im glad u understand.. math isnt my best subject

OpenStudy (anonymous):

i appreciate your help though

OpenStudy (zarkon):

sorry I couldn't be of more assistance.

OpenStudy (anonymous):

your fine.. i understand a lil more than i did before i posted

OpenStudy (zarkon):

good